Please help me solve this equation : x-3/x+2=x+2/x-3
Thank you !

Answer by satyareddy22(84)

x-3/x+2=x+2/x-3
(x-3)^2=(x+2)^2
x^2-6x+9=x^2+4x+4
10x=5
x=5/10=1/2
x=1/2